Operations Report – January 13th


 Operations Report

Crew 219 Operations Report 13-JAN-2020

Sol: 8

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 1.5

From what time last night: 18:00

To what time this morning: *failure stamp 19:30

List of any additional daytime hours when the generator was run:

Start: 06:00

Tuned off: 15:00

Hours run: 9

Solar- SOC% (Before the generator is run at night): 83%

06:30 = 10%; 07:00 = 14%; 08:30 = 22%; 10:00 = 33%; 11:45 = 46%; 12:30
= 51%; 13:30 = 55%; 14:30 = 100%; 15:30 = 98%; 16:30 = 90%; 17:30 =
85%; 18:00 = 83%; 18:30 = 81%;

Diesel Reading - 60%

Station Propane Reading- 39%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147647,2

Water (static tank) (gal) – 384

Static to loft Pump used –Yes

Water in Green Hab (gal): 166.21

Water in Science Dome (gal): 0

Toilet Tank Emptied –Yes

Diemos rover used: Still in the Work shop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

Starting Hours: 124.9

Beginning Charge: 100%

Ending Hours: 125.4

Ending Charge: 27%

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 130.8

Beginning Charge: 100%

Ending Hours: 131.3

Ending Charge: 38%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where?

CrewCar used and why, where?

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Nominal

Summary of Hab: Generator had a power malfunction causing no charge
with the engine running. SOC became under leveled before 06:00 and
safety mode was initiated.

Summary of Science Dome operations: Temperatures are hovering around 4
°C during the day and 0 °C during the night.

Summary of Ram operations: Power converter began a high-pitched
whining noise during the day. Outpost and CO assessed the problem.
Outpost advised to include it in the notes of the operation report.

Summary of any observatory issues: Nominal

Summary of Health and safety issues: See HSO report

Questions, concerns, and requests to Mission Support: None. 
Copyright © The Mars Society. All rights reserved. | Main Site